# An In-depth Evaluation of GPT-4 in Sentence Simplification with Error-based Human Assessment

### v1

@@ 6.1.1. Error Co-occurrence
Multiple types of error may co-occur in a simplification output. Consider the following example where simplification was generated by Control-T5 model:
@@
In this example, Coreference and Altered Meaning-Lexical errors co-occur. The simplification process replaces President Barack Obama with he, leading to a coreference error. Additionally, the phrase the site for his library is oversimplified to a new library thus altering the original meaning.
We found that on average, GPT-4-generated erroneous simplification outputs contained $1.09\pm\scriptstyle{0.28}$ unique errors, while Control-T5-generated ones contained $1.06\pm\scriptstyle{0.25}$. This was calculated by dividing the sum of unique errors in each erroneous simplification output by the total number of these simplification outputs. This suggests that erroneous simplification outputs typically include just one error type on both Control-T5 and GPT-4.
@@ 6.1.2. Distribution of Same Errors
Section 6.1.1 indicates that an erroneous simplification output contains a unique error type on average, then what is the distribution of the same error types in those simplification outputs? The same type of error can occur multiple times within a single simplification output. Consider the following example where the simplification output was generated by GPT-4 model:
@@
In this example, Lack of Simplicity-Lexical error appears twice. The simplification uses more difficult words, sole and permitted, to replace only and allowed, respectively. We plotted the label-wise distribution for both models, as illustrated in Figure 3. This figure shows that in both models, each type of error occurs once in most of the erroneous simplification outputs, and only a small fraction of simplification outputs exhibit the same error type more than once. The maximum repetition of the same error type is capped at three.
Figure 3. Label-wise distribution: Each bar indicates the number of sentences containing a specific type of error, with error frequency ranging from $1$ to $3$.
@@ 6.1.3. Characteristic Errors in Models
We quantitatively assessed the frequency of different error types in simplification outputs generated by GPT-4 and Control-T5. The results, presented in Table 6, indicate differences in error tendencies between the two models.
GPT-4’s errors are predominantly from Lack of Simplicity-Lexical ($100$ occurrences) and Altered Meaning-Lexical ($75$ occurrences), showing its propensity to employ complex lexical expressions or misinterpret meanings through lexical choices. Control-T5, in contrast, displays a broader range of main error types, with notably higher frequencies in Altered Meaning-Lexical ($176$ occurrences), Coreference ($104$ occurrences), and Hallucination ($43$ occurrences). This suggests difficulties with preserving original lexical meanings, ensuring referential clarity, and avoiding irrelevant information for Control-T5. On both models, the occurrences of errors in lexical aspect (Lack of Simplicity-Lexical, Altered Meaning-Lexical, Coreference, Repetition) surpass the occurrences of errors in structural aspect (Lack of Simplicity-Structural, Altered Meaning-Structural) as a general tendency.
Further analysis on Newsela suggests dataset-specific challenges. In comparison to results on Turk and ASSET, Control-T5 generated significantly more Coreference errors than GPT-4. After a manual inspection, we found that a possible reason is the high occurrence of coreference within this dataset and Control-T5 tends to overfit during fine-tuning.
Table 6. Comparison of the number of each error type in GPT-4 and Control-T5 generated simplification outputs
Turk ASSET Newsela Total
GPT-4 T5 GPT-4 T5 GPT-4 T5 GPT-4 T5
Lack of Simplicity - Lexical $21$ 00 $47$ $4$ $32$ 00 $100$ $4$
Lack of Simplicity - Structural 00 $8$ $5$ $7$ $3$ 00 $8$ $15$
Altered Meaning - Structural $2$ $6$ $4$ $7$ $2$ $2$ $8$ $15$
Altered Meaning - Lexical $25$ $79$ $19$ $76$ $31$ $21$ $75$ $176$
Coreference $1$ $7$ $1$ $1$ $12$ $96$ $14$ $104$
Repetition 00 $6$ 00 $1$ 00 00 00 $7$
Hallucination 00 $25$ $4$ $15$ $6$ $3$ $10$ $43$
Total $49$ $131$ $80$ $111$ $86$ $122$ $215$ $364$
@@ 6.2. Likert Scale Rating
In this section, we compared model performances across various dimensions and datasets by averaging annotators’ ratings summarized in Table 7. As a general tendency, GPT-4 again consistently outperforms Control-T5 across all datasets, indicating a preference among annotators for the GPT-4’s simplification quality.
Table 7. Average Ratings
Turk ASSET Newsela
Dimension GPT-4 Control-T5 GPT-4 Control-T5 GPT-4 Control-T5 Reference
Fluency $2.99$ $2.98$ $2.99$ $2.98$ $3.00$ $2.99$ $2.97$
Meaning Preservation $2.90$ $2.43$ $2.88$ $2.41$ $2.66$ $1.16$ $1.67$
Simplicity $2.79$ $2.87$ $2.69$ $2.91$ $2.82$ $2.97$ $2.76$
Total $2.89$ $2.76$ $2.85$ $2.77$ $2.83$ $2.37$ $2.47$
For fluency, both models demonstrate high fluency levels, indicated by the average ratings approach three. This suggests that both GPT-4 and Control-T5 generated grammatically correct simplifications without significant differences in fluency. In terms of meaning preservation, GPT-4 outperforms Control-T5 across all three datasets: achieving scores of $2.90$ versus $2.43$ on Turk, $2.88$ versus $2.41$ on ASSET, and $2.66$ versus $1.16$ on Newsela. Conversely, for simplicity, GPT-4’s ratings are slightly lower than those of Control-T5, though the disparities are less pronounced than those observed for meaning preservation: $2.79$ versus $2.87$ on Turk, $2.69$ versus $2.91$ on ASSET, and $2.82$ versus $2.97$ on Newsela. This contrast suggests that Control-T5 may be slightly better at generating simpler outputs but at the cost of losing a significant portion of the original meaning.
On the Newsela dataset, we observed that Control-T5 often simply deletes long segments of the source sentence, leaving only trivial changes to the remaining parts. Below is an example:
@@
Obama said that children today still fail to meet their full potential, even if there are no mobs with bricks and bottles to ward young black women away from higher education.
Control-T5 Simplification Output::
Even if there are no mobs with bricks and bottles to keep young black women from higher education.
GPT-4 Simplification Output::
Obama said that children aren’t reaching their full potential, even when there are no threats stopping them from getting a better education.
In this example, Control-T5 removed the extensive segment preceding ‘even if…’, resulting in the loss of significant information. In contrast, GPT-4 retained the majority of the information, resulting in higher scores on meaning preservation. The last column of Table 7 shows average ratings for reference simplifications in the Newsela test set, which reveals that Newsela’s references permit to sacrifice of meaning preservation for simplicity. Control-T5 might have adopted a deletion-heavy approach during its training on the Newsela dataset, even heavier than the deletion degree of the Newsela dataset itself. It adversely affects its ability to preserve the original sentence’s meaning.
6.3. Discussion: Other Issues Resolved by LLM
Some Control-T5-generated simplification outputs are reported by our annotators to fail in meeting the satisfaction criteria. These cases include issues such as redundancy, lack of common sense, and inappropriate changes in focus, which were not noted with GPT-4’s simplification outputs. This section provides an enumeration of such cases, alongside examples for each category.
@@ 6.3.1. Redundancy
The simplifications introduced redundancies, failing to contribute meaningfully to the simplification of the sentence or to enhance clarity. In the example below, Control-T5 replaced ‘biochemist’ with ‘biochemist and scientist’, which seems redundant due to the overlapping parts in meanings. In contrast, GPT-4 avoided this issue.
@@
GPT-4 Simplification Output::
Their granddaughter, Hélène Langevin-Joliot, teaches nuclear physics at the University of Paris. Their grandson, Pierre Joliot, named after Pierre Curie, is a famous biochemist.
@@ 6.3.2. Lack of Common Sense
Simplifications that result in logical inconsistencies or nonsensical interpretations. In the example below, Control-T5 illogically suggested that Orton gave birth to his wife, indicating a lack of common sense. In contrast, GPT-4 avoided this issue by correctly suggesting that Orton and his wife welcomed a baby girl, aligning with human common sense understanding.
@@
GPT-4 Simplification Output::
Orton and his wife had a baby girl named Alanna Marie Orton on July 12, 2008.
@@ 6.3.3. Change of Focus
Simplifications that inappropriately alter the original sentence’s focus, leading to misleading interpretations or factual inaccuracies. In the first example below, Control-T5 shifted the focus from the type of piece Opus 57 is to the mere fact that Chopin composed it, while GPT-4 correctly kept the focus of the source sentence. In the second example, Control-T5 inaccurately replaced ‘being proxies for Thaksin’ with ‘Thaksin was a friend of the PAD’, changing the sentence’s meaning substantially. Conversely, GPT-4 did not appear this kind of inference mistake.
Misleading interpretation:
@@
GPT-4 Simplification Output::
Frédéric Chopin’s Opus 57 is a lullaby for a single piano.
Unfactual inference:
@@
The PAD called for the resignation of the governments of Thaksin Shinawatra, Samak Sundaravej, and Somchai Wongsawat, whom the PAD accused of being proxies for Thaksin.
Control-T5 Simplification Output::
The PAD called for the resignation of the governments of Thaksin Shinawatra, Samak Sundaravej, and Somchai Wongsawat. The PAD said that Thaksin was a friend of the PAD.
GPT-4 Simplification Output::
The PAD asked for the resignation of the governments of Thaksin Shinawatra, Samak Sundaravej and Somchai Wongsawat, whom the PAD accused of being representatives for Thaksin.
@@ 7. Meta-Evaluation of Automatic Evaluation Metrics
Due to the high cost and time requirements of human evaluation, automatic metrics are preferred as a means of obtaining for faster and cheaper evaluation of simplification models. Previous studies have explored the extent of widely-used metrics in sentence simplification can assess the quality of outputs generated by neural systems (Sulem et al., 2018a; Alva-Manchego et al., 2021; Tanprasert and Kauchak, 2021). However, it remains uncertain whether these metrics are adequately sensitive and robust to differentiate the quality of simplification outputs generated by advanced LLMs, i.e., GPT-4, especially given the generally high performance. To fill this gap, we performed a meta-evaluation of commonly used automatic metrics in both sentence and corpus levels, utilizing our human evaluation data.
@@ 7.1.1. Sentence-level Metrics
  • BERTScore (Zhang et al., 2020) provides similarity scores (precision, recall, and f1) for each token in the candidate sentence against each token in the reference, leveraging BERT’s (Devlin et al., 2019) contextual embeddings.
@@ 7.1.2. Corpus-level Metrics
  • FKGL (Kincaid et al., 1975) evaluates readability by combining sentence and word lengths. Lower values indicate higher readability. The FKGL score starts from $-3.40$ and has no upper bound.
@@ 7.2. Sentence-Level Results
To assess sentence-level metrics’ ability on differentiating the sentence-level simplification quality, we explore the correlation between those metrics and human evaluations by employing the point-biserial correlation coefficient (Glass and Hopkins, 1995; Linacre, 2008), utilizing the scipy package (Virtanen et al., 2020) for calculation [^7] The point-biserial correlation coefficient was chosen because our human labels are mostly binary while evaluation metric scores are continuous.. This coefficient ranges from $-1$ and $+1$, where 00 indicates no correlation.
@@
Given the data imbalance between sentences with and without errors, and between high-quality and low-quality simplification, we report our findings using both raw data and downsampled (DS) data to balance the number of class samples.
@@ 7.2.1. Identification of the presence of errors
For all $1800$ simplification outputs in Task $1$, each simplification output was classified as containing errors (labeled as $1$) or no error (labeled as 00). We then computed the correlation coefficients between these labels and the metric scores. The results, presented in Table 8, indicate that none of the metrics effectively identify erroneous simplifications, as evidenced by point-biserial correlation coefficients being near zero.
@@
All GPT-4 Control-T5
Raw DS ($511$) Raw DS ($182$) Raw DS ($329$)
LENS $-0.04$ $-0.03$ $-0.04$ $-0.04$ $-0.13$ $-0.11$
BERT precision $-0.07$ $-0.11$ $-0.10$ $-0.16$ $-0.05$ $-0.03$
BERT recall $-0.12$ $-0.16$ $-0.14$ $-0.20$ $-0.12$ $-0.10$
BERT f1 $-0.09$ $-0.13$ $-0.11$ $-0.18$ $-0.07$ $-0.05$

### v2

@@ 6.1.1. Error Co-occurrence
Multiple types of errors may co-occur in a simplification output. Consider the following example where simplification is generated by the Control-T5 model:
@@
In this example, Coreference and Altered Meaning-Lexical errors co-occur. The simplification process replaces President Barack Obama with he, leading to a coreference error. Additionally, the phrase the site for his library is oversimplified to a new library, thus altering the original meaning. We find that on average, GPT-4-generated erroneous simplification outputs contain $1.08\pm\scriptstyle{0.28}$ unique errors, Qwen2.5-72B-generated ones contain $1.04\pm\scriptstyle{0.22}$, Llama-3.2-3B-generated ones contain $1.14\pm\scriptstyle{0.4}$, and Control-T5-generated ones contain $1.05\pm\scriptstyle{0.24}$. These averages are calculated by dividing the sum of unique errors in each erroneous simplification output by the total number of these simplification outputs. This indicates that erroneous simplification outputs across all models typically include only one type of error.
@@ 6.1.2. Distribution of Same Errors
Section 6.1.1 indicates that an erroneous simplification output contains a unique error type on average, then what is the distribution of the same error types in those simplification outputs? The same type of error can occur multiple times within a single simplification output. Consider the following example where the simplification output is generated by GPT-4:
@@
In this example, Lack of Simplicity-Lexical error appears twice. The simplification uses more difficult words, sole and permitted, to replace only and allowed, respectively. Upon a close observation, we find that across all models, each type of error occurs once in most of the erroneous simplification outputs, and only a small fraction of simplification outputs exhibit the same error type more than once. The maximum repetition of the same error type is capped at four.
@@ 6.1.3. Characteristic Errors in Models
We quantitatively analyze the frequency of different error types in simplification outputs generated by models. The results, shown in Figure 4, indicate variations in error tendencies across models.
LLMs Outperform Control-T5
Consistent with the sentence-level results in Figure 3, Control-T5 generates more errors overall ($350$ occurrences) than the LLM group ($211$ for GPT-4, $172$ for Qwen2.5-72B, and $326$ for Llama-3.2-3B after excluding SimPA). Among the LLMs, Qwen2.5-72B produces the fewest errors, followed by GPT-4, with Llama-3.2-3B generating significantly more errors ($202$ for Qwen2.5-72B, $285$ for GPT-4, and $405$ for Llama-3.2-3B). Notably, Qwen2.5-72B performs best in four out of seven error categories, suggesting that while larger LLMs generally perform better, performance may not always scale directly with model size in simplification.
Lexical Paraphrasing is the Biggest Challenge
Both GPT-4 and Qwen2.5-72B show similar tendencies, with errors predominantly from Lack of Simplicity-Lexical ($144$ for GPT-4 and $98$ for Qwen2.5-72B) and Altered Meaning-Lexical ($94$ for GPT-4 and $59$ for Qwen2.5-72B). This reflects their propensity to employ complex lexical expressions or misinterpret meanings through lexical choices, though Qwen2.5-72B performs better in these categories. Control-T5 shows notably high frequencies in Altered Meaning-Lexical ($176$ occurrences) and Coreference ($104$ occurrences). This indicates difficulties with preserving original lexical meanings and ensuring referential clarity. Across all models, errors in lexical aspect (Lack of Simplicity-Lexical, Altered Meaning-Lexical, Coreference, Repetition) surpass the occurrences of errors in structural aspect (Lack of Simplicity-Structural, Altered Meaning-Structural) as a general tendency.
Newsela Poses Coreference Resolution Challenge
Further analysis of Newsela reveals dataset-specific challenges. Compared to Turk and ASSET, Control-T5 generates significantly more Coreference errors ($7$ for Turk, $1$ for ASSET, and $96$ for Newsela). After a manual inspection, we find that a possible reason is the high occurrence of coreference within this dataset, and Control-T5 tends to overfit during fine-tuning.
Llama-3 is Prone to Repetition Error
Remarkably, for Llama-3.2-3B, while paraphrasing remains a significant issue, errors such as Repetition and Hallucination are notably more frequent than in other models. For repetition, some of the errors may stem from the model’s misunderstanding of the prompt crafted for Newsela, obtaining from prompt engineering on GPT-4 (see Section 4.1.2). That is, providing $3$-shot examples with multiple simplification references under each example. Llama-3.2-3B appears to combine multiple simplifications into a single output, leading to repetitive content. Below is an example:
Source::
But landowner Gene Pfeifer refused to give up his 3-acre riverfront property in the middle of the proposed library site.
Llama::
Gene Pfeifer didn’t want to sell his 3-acre land. Gene Pfeifer refused to sell his land. Gene Pfeifer didn’t want to give up his 3-acre property.
Figure 4. Error type distribution across models: each bar represents the count of a specific error type for each model. To ensure a fair comparison between LLMs and Control-T5, we use ‘/’ and ‘.’ in the graph to distinguish results before (‘T&A&N’) and after incorporating SimpA.
@@ 6.2. Likert Scale Rating
In this section, we compare model performances across various dimensions and datasets by averaging annotators’ ratings. Results show that, as a general tendency, LLMs again consistently outperform Control-T5 across all datasets, indicating a preference among annotators for the LLMs’ simplification quality. Among the LLM group, GPT-4 and Qwen2.5-72B show comparable performance, consistently rated higher than Llama-3.2-3B across all datasets.
Figure 5. Comparison of meaning Preservation and simplicity across models and datasets
For fluency, all models demonstrate high fluency levels, indicated by the average ratings approaching three. This suggests that these models generate grammatically correct simplifications without significant differences in fluency. In terms of meaning preservation, GPT-4 and Qwen2.5-72B outperform Llama-3.2-3B and Control-T5 across all datasets: achieving scores of ($2.90$ (GPT-4), $2.91$ (Qwen)) versus ($2.47$ (Llama), $2.43$ (T5)) on Turk, ($2.88$, $2.92$) versus ($2.70$, $2.41$) on ASSET, ($2.66$, $2.62$) versus ($1.78$, $1.16$) on Newsela, and ($2.89$, $2.90$) versus ($2.43$) on SimPA. Conversely, for simplicity, GPT-4 and Qwen2.5-72B’s ratings are comparable or slightly lower than those of Llama-3.2-3B and Control-T5, though the disparities are less pronounced than those observed for meaning preservation: ($2.79$ (GPT-4), $2.94$ (Qwen)) versus ($2.90$ (Llama), $2.87$ (T5)) on Turk, ($2.69$, $2.86$) versus ($2.76$, $2.91$) on ASSET, ($2.82$, $2.93$) versus ($2.68$, $2.97$) on Newsela, and ($2.92$, $2.94$) versus ($2.94$) on SimPA. To facilitate a clearer comparison, we provide Figure 5, which visually contrasts the models’ performance in meaning preservation and simplicity. The contrast suggests that Llama-3.2-3B and Control-T5 may be comparably good or slightly better at generating simpler outputs but at the cost of losing a significant portion of the original meaning.
On the Newsela dataset, we observe that Control-T5 often simply deletes long segments of the source sentence, leaving only trivial changes to the remaining parts. Below is an example:
@@
Sun said his fascination with the past started with his grandfather, who taught him the old accents and ways of Beijing, including Beijing opera.
Control-T5::
Sun said his fascination with the past started with his grandfather.
GPT-4::
Sun’s grandfather taught him about the culture of old Beijing, which sparked his interest in the past.
Qwen::
Sun’s interest in the past began with his grandfather, who taught him about old Beijing, including its opera.
Llama::
Sun’s interest in the past started with his grandfather, who taught him about Beijing’s old ways.
In this example, the source sentence consists of a main clause and a relative clause. Control-T5 removes the relative clause entirely while leaving the main clause unchanged, leading to significant information loss and limited transformation diversity. In contrast, GPT-4, Qwen2.5-72B, and Llama-3.2-3B tend to retain information from both parts, achieving higher scores for meaning preservation. The average rating for reference simplifications in the Newsela test set is $1.67$, which reveals that Newsela’s references sacrifice meaning preservation for simplicity. Control-T5 might have adopted a deletion-heavy approach during its training on the Newsela dataset, even heavier than the deletion degree of the Newsela dataset itself. It adversely affects its ability to preserve the original sentence’s meaning.
6.3. Discussion: Additional Observations
During the annotation process, we observed several nuanced phenomena that were difficult to fit into specific error categories. Some cases fail to meet the criteria for satisfactory simplifications, including redundancy, lack of common sense, and change of focus. Other cases could be controversial, including the addition of factual information not inferable from the source sentence. This section outlines the models where these phenomena were observed and provides the number of reported cases during error identification annotation, along with examples for each category.
@@ 6.3.1. Redundancy
Simplifications introduced redundancies, failing to contribute meaningfully to the simplification of the sentence or to enhance clarity. This was only observed in Control-T5’s simplifications with four reported cases. In the example below, Control-T5 replaces ‘biochemist’ with ‘biochemist and scientist, ’ which seems redundant due to the overlapping parts in meanings.
@@ 6.3.2. Lack of Common Sense
Simplifications that result in logical inconsistencies or nonsensical interpretations. This was only observed in Control-T5’s simplifications with three reported cases. In the example below, Control-T5 illogically suggests that Orton gave birth to his wife, indicating a lack of common sense.
@@ 6.3.3. Change of Focus
Simplifications that inappropriately alter the original sentence’s focus, leading to misleading interpretations. This was only observed in Control-T5 with four reported cases and Llama-3.2-3B with six. In the first example below, Control-T5 shifts the focus from the type of piece Opus 57 to the mere fact that Chopin composed it. Similarly, in the second example, Llama-3.2-3B redirects attention from the agreement and actions of other judges to the federal court’s decision itself.
@@
Other judges agreed with the federal court’s decision and started marrying same-sex couples in the morning.
Llama::
The federal court ruled that same-sex couples could get married.
6.3.4. Factual Information Not Inferable from the Source Sentence
We found cases where information not explicitly present in the source sentence was added to the simplifications. This was observed in all models, with four reported cases in GPT-4, $12$ in Qwen2.5-72B, five in Llama-3.2-3B, and $12$ in Control-T5. These additions are generally factual and, although not inferable from the source sentence, were verified to be factual using online sources. This type of information can be controversial as it does not strictly adhere to the input. However, it may facilitate the reader’s understanding of the source sentence. We did not classify these cases as errors but instead documented them along with external resources, such as website links, to verify their accuracy. For example, in the case below, “Lincoln’s assassination” cannot be inferred directly from the source sentence. However, Qwen2.5-72B includes this detail, likely drawing on its internal knowledge by linking the provided date and named entities. In such situations, we verified the added information online. If the information was factual, we did not classify it as an error.
  • Source: For example, there’s a letter of sympathy from Queen Victoria to Mary Todd Lincoln on April 29, 1865, calling his assassination “so terrible a calamity”.
  • Qwen: Queen Victoria wrote a letter of sympathy to Mary Todd Lincoln about Lincoln’s assassination.
@@ 7. Meta-Evaluation of Automatic Evaluation Metrics
Due to the high cost and time requirements of human evaluation, automatic metrics are preferred as a means of obtaining faster and cheaper evaluation of simplification models. Previous studies have explored the extent of widely-used metrics in sentence simplification can assess the quality of outputs generated by neural systems (Sulem et al., 2018a; Alva-Manchego et al., 2021; Tanprasert and Kauchak, 2021). However, it remains uncertain whether these metrics are adequately sensitive and robust to differentiate the quality of simplification outputs generated by advanced LLMs, i.e., GPT-4, especially given the generally high performance. To fill this gap, we perform a meta-evaluation of commonly used automatic metrics at both sentence and corpus levels, utilizing our human evaluation data.
@@ 7.1.1. Sentence-level Metrics
  • BERTScore (Zhang et al., 2020) provides similarity scores (precision, recall, and f1) for each token in the candidate sentence against each token in the reference, leveraging BERT’s (Devlin et al., 2019) contextual embeddings. In this study, we use the f1-score as we observed that the trends of recall, precision, and f1 are similar.
@@ 7.1.2. Corpus-level Metrics
  • FKGL (J. Peter et al., 1975) evaluates readability by combining sentence and word lengths. Lower values indicate higher readability. The FKGL score starts from $-3.40$ and has no upper bound.
@@ 7.2. Sentence-Level Results
To assess sentence-level metrics’ ability on differentiating the sentence-level simplification quality, we explore the correlation between those metrics and human evaluations by employing the point-biserial correlation coefficient (Glass and Hopkins, 1995; Linacre, 2008), utilizing the scipy package (Virtanen, 2020) for calculation [^8] The point-biserial correlation coefficient was chosen because our human labels are mostly binary while evaluation metric scores are continuous.. This coefficient ranges from $-1$ and $+1$, where 00 indicates no correlation.
@@
Given the data imbalance between sentences with and without errors and between high-quality and low-quality simplification, we report our findings using both raw data and downsampled (DS) data to balance the number of class samples.
@@ 7.2.1. Identification of the Presence of Errors
For all $4,500$ simplification outputs in Task $1$, each simplification output is classified as containing errors (labeled as $1$) or no error (labeled as 00). We then compute the correlation coefficients between these labels and the metric scores. The results, presented in Table 5, indicate that none of the metrics effectively identify erroneous simplifications, as evidenced by point-biserial correlation coefficients being near zero.
@@
All GPT-4 Qwen2.5-72B Llama-3.2-3B Control-T5
Raw DS ($1073$) Raw DS ($240$) Raw DS ($182$) Raw DS ($332$) Raw DS ($319$)
LENS $-0.16$ $-0.18$ $-0.10$ $-0.16$ $-0.10$ $-0.20$ $-0.25$ $-0.26$ $-0.14$ $-0.14$
BERT f1 $-0.12$ $-0.12$ $-0.12$ $-0.14$ $-0.03$ $-0.04$ $-0.20$ $-0.25$ $-0.12$ $-0.12$